OpenStudy (anonymous):

2. Write the system of equations represented by the matrix. −1 1 2 1−67

OpenStudy (anonymous):

is that supposed to be one row?

OpenStudy (anonymous):

I think so, that's all it says

OpenStudy (anonymous):

This matrix? $$ \begin{bmatrix} −1& 1& ~2 & 1&−6&7 \end{bmatrix} $$

OpenStudy (anonymous):

It might be :( My teacher just sent the document like that. Can it be done?

OpenStudy (anonymous):

can i see the document

OpenStudy (anonymous):

This might be what they want, i'm not sure $$\begin{bmatrix} −1& 1& ~2 & 1&−67 \end{bmatrix} \iff -1x + 1y + 2z + 1w = -67 $$

OpenStudy (anonymous):

Ok thank you!! I appreciate it